OpenStudy (anonymous):

4 over 1-x= 7 over 1+x

OpenStudy (anonymous):

4/(1-x) = 7/(1+x) Correct?

OpenStudy (anonymous):

4(1+x)=7(1-x) ?

OpenStudy (anonymous):

What are you trying to prove here?

OpenStudy (anonymous):

Are you trying to solve for x??

OpenStudy (anonymous):

im trying to find the solution yes im trying to find x

OpenStudy (anonymous):

Ok are you working from a textbook? If you are what is the answer??

OpenStudy (anonymous):

no its a homework problem. i know i'm supose to cross mutiply which is would be 4(1+x)=7(1-x) but i'm not sure what to do after that

OpenStudy (anonymous):

Ok so after that we can distribute the 4 and the 7 into the binomial. So do that and show me your answer

OpenStudy (anonymous):

i dont know what ya mean. am i suppose to do this 4+4x=7-7x

OpenStudy (anonymous):

Yes :) Now collect like terms :)

OpenStudy (anonymous):

So what did you get?

OpenStudy (anonymous):

-3/11?

OpenStudy (anonymous):

or 11/3

OpenStudy (anonymous):

-3/11 is correct :)

OpenStudy (anonymous):

Congratulations :) Did I help in anyway??

OpenStudy (anonymous):

yes thankyou

OpenStudy (anonymous):

Yay :)

OpenStudy (anonymous):

double check that... i think it should be positive 3/11

OpenStudy (anonymous):

Nope I checked :P

OpenStudy (anonymous):

@dpaInc I want to know your reason

OpenStudy (anonymous):

I know why @Shauna8 cross multiplied wrong

OpenStudy (anonymous):

@Shauna8 Your cross multiplication should say \[7(1+x) = 4(1-x)\] Not \[4(1+x) = 7(1-x)\] My mistake :P

OpenStudy (anonymous):

\(\large 4+4x=7-7x \) \(\large 4+4x\color {red}{-4+7x}=7-7x\color {red}{-4+7x} \) add 7x and subtract 4 to both sides \(\large 11x=3 \) solve from here...

OpenStudy (anonymous):

so its 3/11

OpenStudy (anonymous):

No it's -3/11